# Backreaction from quantum fluxes at the Kerr inner horizon

### v1

@@ IV Backreaction
Analyses of semiclassical backreactions carry with them intricate subtleties. Ideally, one would like to find a solution to the $\o$ller-Rosenfeld semiclassical field equations, Eq. (1), to see how the geometry and the quantum field coevolve over time. However, as can be gleaned from the calculations of the previous sections, the RSET is difficult to calculate and currently is only known for a select handful of highly symmetric vacuum spacetimes. Thus, the present calculations will apply only to a weak backreaction domain, where one is still far enough from the inner horizon that the geometry can be well-approximated by the Kerr metric but close enough that the semiclassical backreaction, of order $(\text{e}^{\kappa_{-}v}\ m_{P}/M)^{2}$ (where $m_{P}$ denotes the Planck mass), is not negligible.
@@ IV.1 Spherical symmetry
However, if instead the radial coordinate $r$ is gauge-fixed to match that of the vacuum RSET and the remaining two metric degrees of freedom are encoded by, for example, the positions of the black hole’s inner and outer horizons, the same field equations then imply dynamical behavior in the horizon structure rather than in the local area element. In particular, the outer horizon will evaporate inward slowly over time, as per Hawking’s famous result (as long as ${\langle T^{+}_{vv}\rangle<0}$, which is always true in the 2D case) [29, 30, 31], while the inner horizon will evaporate rapidly outward (inward) when ${\langle T^{-}_{vv}\rangle}$ is negative (positive). The inner horizon in this case will move extremely quickly, at a timescale on the order of the black hole’s light-crossing time [31].
@@ IV.2 Axisymmetry: initial tendencies
Now consider the case of black holes with non-zero rotation. Here it is clear that the outcome must be more complicated than in the spherically symmetric case, since the local interior geometry will no longer uniformly contract or expand across the entire inner horizon; instead, ${\langle T^{-}_{vv}\rangle}$ is both positive and negative across different latitudes of the same black hole (see Fig. 5). An excellent backreaction analysis for the case of Kerr-de Sitter black holes was recently [17] carried out by decomposing the spacetime into a set of double-null hypersurfaces gauge-fixed to match the null coordinates in ${\langle T^{-}_{\mu\nu}\rangle}$. [^4] The fixing comes in the identification of the affine parameter $\lambda$, taken to be constant along each hypersurface, with the Kruskal coordinate $V_{-}\propto\text{e}^{-\kappa_{-}v}$. Along each hypersurface, the induced metric can be written in a completely general form as
@@
for the area element $\gamma^{2}$ and the arbitrary functions $\alpha$ and $\tau$ that physically relate to notions of shear and twisting of null geodesics as the hypersurface is moved toward the would-be inner horizon at ${\lambda\to 0}$, or equivalently, ${v\to\infty}$. The semiclassical Einstein equations and Raychaudhuri equations can then be analyzed to yield
@@
where $\beta$ measures the amount of twisting in the $\varphi$-direction for light rays perpendicular to the surfaces of constant $\lambda$ as ${\lambda\to 0}$.
From Eq. (12a) it can be gleaned that for positive (negative) ${\langle T^{-}_{vv}\rangle}$, the divergent RSET component ${\langle T^{-}_{\lambda\lambda}\rangle\sim\lambda^{-2}}$ will also be positive (negative) and will cause $\gamma$ to contract (expand) as ${\lambda\to 0}$. In the inevitable intermediate cases where $\gamma$ neither contracts nor expands, Eq. (12b) then predicts that a non-zero $\langle T^{-}_{v\varphi}\rangle$ will cause an infinite local twisting. Thus, one reproduces the same behaviors predicted in the spherically symmetric case (i.e., the local blow-up or shrinking of the geometry near the inner horizon dictated by the sign of ${\langle T^{-}_{vv}\rangle}$), with the addendum that even in the cases where the geometry neither contracts nor expands, a subdominant component of the RSET will still generically cause the local geometry to diverge in a shearing manner.
Two comments concerning the above analysis are worth mentioning. First, while the backreaction is performed with sufficient generality to reproduce the effects predicted in the spherical case and predict additional rotational effects, the restriction of the metric to null hypersurface cuts of constant $v$ does not allow for a full global analysis of the spacetime’s evolution and does not encompass the most general axisymmetric geometry possible (it for example washes out any dynamical information related to the outgoing $u$ coordinate and its coupling with the other degrees of freedom). But more importantly, the analysis does not invite any obvious path forward to understand how the spacetime evolves beyond these initial tendencies. Does the semiclassical inner-horizon instability remain confined to a (meta)stable interior singularity (or even a regular configuration), or will inflationary perturbations spread like a wildfire until they destroy the black hole from the inside out? To address these problems at least partially, consider the complementary analysis below.
@@
The properties of the semiclassical radiation produced from a quantum scalar field in the vacuum Kerr spacetime already seem to rule out a number of Kerr generalizations. For example, the 3-function class of metrics put forward in Ref. [35] claims to encompass a majority of the currently-proposed stable endpoints of black hole evolution (aside from complete evaporation or gravastar-like objects); however, the Einstein tensor for these metrics always contains a vanishing $rt$-component, despite the fact that the semiclassical RSET calculated here possesses a non-vanishing (and in fact diverging) $rt$-component ${\langle T^{-}_{rt}\rangle}={\left(\langle T^{-}_{vv}\rangle-\langle T^{-}_{uu}\rangle\right)R^{-2}\Delta_{r}^{-1}}$. Presumably such black hole solutions that are regular and instability-free cannot form from an initial Kerr-like collapse.
@@ IV.3 Axisymmetry: mass inflationary approach
    The spacetime is initialized with a Kerr vacuum geometry, and the quantum scalar field is initialized with an Unruh vacuum state mimicking the effects of a gravitational collapse sufficiently far in the past.
@@
    The spacetime maintains conformal Hamilton-Jacobi separability [5]; i.e., the equations of motion for massless particles are separable in spheroidal coordinates (but not necessarily for massive particles, as in the case of strict separability seen for Kerr).
Unfortunately, none of these assumptions are actually true for astrophysically realistic black holes. The accretion flows observed by the GRAVITY instrument [37] and the Event Horizon Telescope Collaboration [38] feature asymmetries in $\varphi$, though in the stationary limit, rigidity theorems suggest that any non-axisymmetric perturbations are part of the “hair” that a black hole will eventually shed [39, 40]. In regard to Assumption (ii), the lack of asymptotic flatness in the Universe has been measured with great precision [41], but any effects from a cosmological horizon on the inner-horizon instability are expected to be negligible, both due to the similarity in the RSET calculations here to those of the Kerr-de Sitter case [17], and because the cosmological constant is currently measured to be $\Lambda\sim 10^{-46}\ M_{\astrosun}^{-2}$ in geometrized units, much too small to have any practical influence. As for Assumption (iii), black holes (and therefore near-inner-horizon geometries) form under chaotic, rapidly evolving conditions, and even after settling down, the most quiescent black holes are still non-vacuum and accrete more than enough radiation to trigger the mass inflation instability and destroy any vacuum inner horizon [8]. Additionally, physically relevant fields in the Standard Model are those with spin 1 (electromagnetic), 1/2 (fermionic), or 2 (gravitational), so using a massless scalar field with spin 0 is a simplification seen as a proxy for, e.g., a single photonic degree of freedom. Even regarding Assumption (iv), which provides a more general notion than the strict separability that itself is justified by the observation of long-lived accretion disks (lack of separability implies chaotic, destabilizing particle orbits [35]), it will be seen in what follows that conformal separability, while valid in the weak backreaction domain, eventually breaks down once the vierbein of Eq. (16) ceases to be diagonal.
@@
The Einstein tensor corresponding to this spacetime is far too intricate to be presented here in any complete, meaningful way. However, at least in the quasi-stationary limit (in particular, if the redefined conformal piece $\rho$ is written as $\rho(r,\theta,t)=\text{e}^{\tilde{v}t}\sqrt{r^{2}+a^{2}\cos^{2}\!\theta}$ for some vanishing small accretion parameter $\tilde{v}$), the field equations have already been analyzed in detail in Refs. [4, 5, 6]. The resulting behavior provides classical justification for the decomposition of Eq. (14), since near the inner horizon, the functions $\Delta_{\theta}$, $\Omega_{r}$, and $\Omega_{\theta}$ remain approximately fixed at their Kerr values near the inner horizon, while the horizon function $\Delta_{r}$ and the inflationary function $\xi$ become the dominant contributors to the evolution of the Einstein tensor. The resulting tensor behaves at least initially like a perfect fluid with equal ingoing and outgoing streams, with dominant inflating tetrad-frame components $G_{00}\sim G_{11}\sim G_{01}\sim\Delta_{r}^{-1}$, where the off-diagonal component $G_{01}$ is suppressed by a factor of $\tilde{v}$. Classically, the evolution can then be continued to show that $\Delta_{r}$ eventually stalls out at some tiny value [5], after which $\xi$ grows large and the spacetime undergoes a series of BKL-like bounces toward a strong, spacelike singularity [7].
How might semiclassical effects modify this picture? The tetrad-frame components of the RSET can be re-expressed in terms of the double-null Eddington-Finkelstein coordinates used throughout this paper:
@@
Far enough away from the inner horizon, all components of the RSET should be completely negligible, as they are suppressed by the Planck scale. Thus, everywhere outside the black hole as well as inside when $v$ is not large, a vacuum source should recover the standard Kerr solution with $\xi=0$ and $\Delta_{r}=(r^{2}+a^{2}-2Mr)/(r^{2}+a^{2})^{2}$. However, in the weak backreaction domain, once the null components of the RSET cease to be vanishingly small, Eqs. (17) indicate that the radial- and time-components of the tetrad-frame Einstein tensor will begin to diverge as $\Delta_{r}^{-1}$ as one takes $\Delta_{r}\to 0$.
Though the RSET components ${\langle T^{-}_{uv}\rangle}$, ${\langle T^{-}_{u\varphi}\rangle}$, and ${\langle T^{-}_{\varphi\varphi}\rangle}$ have not been computed explicitly, one can make the assumption that their contributions to Eqs. (17) will be subdominant compared to that of the double-null components, and more importantly, that they will not contrive to cause any of the specific combinations in the equations above to cancel exactly. Even if they do, then the classical mass inflation phenomenon described in Refs. [4, 5, 6] will take over the evolution and a spacelike singularity will form. But based on the numerical results in Sec. III, it is apparent that ${\langle T^{-}_{00}\rangle}$ and ${\langle T^{-}_{11}\rangle}$ are the dominant contributors to the evolution in the weak backreaction domain, while the off-diagonal components ${\langle T^{-}_{01}\rangle}$ are also important but initially much smaller than their diagonal counterparts (recall from Fig. 3 that the difference ${\langle T^{-}_{uu}\rangle}-{\langle T^{-}_{vv}\rangle}$ is almost always several orders of magnitude smaller than either individual component).
@@
$-2\partial_{r_{*}}\ln\left(\frac{\rho}{1-\Omega_{r}\Omega_{\theta}}\right)\partial_{r_{*}}\xi\bigg{)}+F_{rt},$ (18a)
where the tortoise coordinate $r_{*}$ is defined by $dr/dr_{*}=\text{e}^{3\xi}R^{2}\Delta_{r}$, and the function $F_{rt}$ encompasses subdominant terms related to the precise nature of any classical accretion contributing to non-stationary conformal dynamics.
The dominant term in Eq. (18) is the one involving the second derivative of the inflationary exponent $\xi$ [5]. Equating this term to the semiclassical source from Eq. (17), which behaves as $\Delta_{r}^{-1}$ near the inner horizon, yields after integration the approximate solution
@@
$-2\Omega_{r}(\langle T^{-}_{u\varphi}\rangle-\langle T^{-}_{v\varphi}\rangle)+\Omega_{r}^{2}\langle T^{-}_{\varphi\varphi}\rangle\Big{)},$ (19a)
as long as $\xi$ remains smaller than its derivatives. In the regime where the double-null components of the RSET dominate over the shearing components, Eq. (19) thus predicts that when the sum ${\langle T^{-}_{uu}\rangle+\langle T^{-}_{vv}\rangle}$ is positive (negative), the spacetime’s conformal factor $\text{e}^{-\xi}$ will abruptly contract (expand) as the inner horizon is approached at $\Delta_{r}\to 0_{-}$. These initial tendencies align precisely with that of the spherically symmetric case, since ${\langle T^{-}_{uu}\rangle+\langle T^{-}_{vv}\rangle}$ is usually the same sign as ${\langle T^{-}_{vv}\rangle}$.
One can continue the evolution of the spacetime by examining the behavior of the dynamical vierbein of Eqs. (16a)–(16d). Under the assumption that the fixed vierbein of Eqs. (15a)–(15d) remains stable as the initial backreaction ignites the inflationary tendencies described above, the geometry will be governed solely by a diagonal, dynamical, homogeneous line element sourced by inflating streams of semiclassical matter. The classical counterpart of this analysis has already been carried out in detail in Ref. [8], yielding the so-called inflationary Kasner metric: as one might anticipate, the streams of matter are amplified in energy density as $\Delta_{r}$ plunges to zero and $\xi$ increases rapidly. Then, the spacetime undergoes a bounce as $\Delta_{r}$ freezes out at a small value and the $\text{e}^{\xi/2}$ term in Eq. (16b) takes over and causes a radial expansion. The solution works for both positive and negative sources of stress-energy, generically producing a series of inflating Kasner-like bounces toward a spacelike singularity.

### v2

@@ IV Backreaction
Analyses of semiclassical backreactions carry with them intricate subtleties. Ideally, one would like to find a solution to the Møller-Rosenfeld semiclassical field equations, Eq. (1), to see how the geometry and the quantum field coevolve over time. However, as can be gleaned from the calculations of the previous sections, the RSET is difficult to calculate and currently is only known for a select handful of highly symmetric vacuum spacetimes. Thus, the present calculations will apply only to a weak backreaction domain, where one is still far enough from the inner horizon that the geometry can be well approximated by the Kerr metric, but close enough that the semiclassical backreaction, of order $(\text{e}^{\kappa_{-}v}\ m_{P}/M)^{2}$ (where $m_{P}$ denotes the Planck mass), is not negligible.
@@ IV.1 Spherical symmetry
However, if instead the radial coordinate $r$ is gauge-fixed to match that of the vacuum RSET and the remaining two metric degrees of freedom are encoded by—for example—the positions of the black hole’s inner and outer horizons, the same field equations then imply dynamical behavior in the horizon structure rather than in the local area element. In particular, the outer horizon will evaporate inward slowly over time, as per Hawking’s famous result (as long as ${\langle T^{+}_{vv}\rangle<0}$, which is always true in the 2D case) [29, 30, 31], while the inner horizon will evaporate rapidly outward (inward) when ${\langle T^{-}_{vv}\rangle}$ is negative (positive). The inner horizon in this case will move extremely quickly, at a timescale on the order of the black hole’s light-crossing time [31].
@@ IV.2 Axisymmetry: Initial tendencies
Now, consider the case of black holes with nonzero rotation. Here it is clear that the outcome must be more complicated than in the spherically symmetric case, since the local interior geometry will no longer uniformly contract or expand across the entire inner horizon; instead, ${\langle T^{-}_{vv}\rangle}$ is both positive and negative across different latitudes of the same black hole (see Fig. 5). An excellent backreaction analysis for the case of Kerr–de Sitter black holes was recently [17] carried out by decomposing the spacetime into a set of double-null hypersurfaces gauge-fixed [^4] The fixing comes in the identification of the affine parameter $\lambda$, taken to be constant along each hypersurface, with the Kruskal coordinate $V_{-}\propto\text{e}^{-\kappa_{-}v}$. to match the null coordinates in ${\langle T^{-}_{\mu\nu}\rangle}$. Along each hypersurface, the induced metric can be written in a completely general form as
@@
for the area element $\gamma^{2}$ and the arbitrary functions $\alpha$ and $\tau$ that physically relate to notions of shear and twisting of null geodesics, as the hypersurface is moved toward the would-be inner horizon at ${\lambda\to 0}$, or equivalently, ${v\to\infty}$. The semiclassical Einstein equations and Raychaudhuri equations can then be analyzed to yield
@@
where $\beta$ measures the amount of twisting in the $\varphi$ direction for light rays perpendicular to the surfaces of constant $\lambda$ as ${\lambda\to 0}$.
From Eq. (12a), it can be gleaned that for positive (negative) ${\langle T^{-}_{vv}\rangle}$, the divergent RSET component ${\langle T^{-}_{\lambda\lambda}\rangle\sim\lambda^{-2}}$ will also be positive (negative) and will cause $\gamma$ to contract (expand) as ${\lambda\to 0}$. In the inevitable intermediate cases where $\gamma$ neither contracts nor expands, Eq. (12b) then predicts that a nonzero $\langle T^{-}_{v\varphi}\rangle$ will cause an infinite local twisting. Thus, one reproduces the same behaviors predicted in the spherically symmetric case (i.e., the local blow-up or shrinking of the geometry near the inner horizon dictated by the sign of ${\langle T^{-}_{vv}\rangle}$), with the addendum that even in the cases where the geometry neither contracts nor expands, a subdominant component of the RSET will still generically cause the local geometry to diverge in a shearing manner.
Two comments concerning the above analysis are worth mentioning. First, while the backreaction is performed with sufficient generality to reproduce the effects predicted in the spherical case and predict additional rotational effects, the restriction of the metric to null hypersurface cuts of constant $v$ does not allow for a full global analysis of the spacetime’s evolution and does not encompass the most general axisymmetric geometry possible (it, for example, washes out any dynamical information related to the outgoing $u$ coordinate and its coupling with the other degrees of freedom). But more importantly, the analysis does not invite any obvious path forward to understand how the spacetime evolves beyond these initial tendencies. Does the semiclassical inner-horizon instability remain confined to a (meta)stable interior singularity (or even a regular configuration), or will inflationary perturbations spread like a wildfire until they destroy the black hole from the inside out? To address these problems at least partially, consider the complementary analysis below.
@@
The properties of the semiclassical radiation produced from a quantum scalar field in the vacuum Kerr spacetime already seem to rule out a number of Kerr generalizations. For example, the 3-function class of metrics put forward in Ref. [35] claims to encompass a majority of the currently proposed stable end points of black hole evolution (aside from complete evaporation or gravastar-like objects); however, the Einstein tensor for these metrics always contains a vanishing $rt$-component, despite the fact that the semiclassical RSET calculated here possesses a nonvanishing (and in fact diverging) $rt$-component ${\langle T^{-}_{rt}\rangle}={\left(\langle T^{-}_{vv}\rangle-\langle T^{-}_{uu}\rangle\right)R^{-2}\Delta_{r}^{-1}}$. Presumably, such black hole solutions that are regular and instability-free cannot form from an initial Kerr-like collapse.
@@ IV.3 Axisymmetry: Mass inflationary approach
    The black hole is initially isolated as in the Kerr vacuum geometry, and the quantum scalar field is initialized with an Unruh vacuum state mimicking the effects of a gravitational collapse sufficiently far in the past.
@@
    The spacetime maintains conformal Hamilton-Jacobi separability [6]—i.e., the equations of motion for massless particles are separable in spheroidal coordinates (but not necessarily for massive particles, as in the case of strict separability seen for Kerr).
Unfortunately, none of these assumptions are actually true for astrophysically realistic black holes. The accretion flows observed by the GRAVITY instrument [37] and the Event Horizon Telescope Collaboration [38] feature asymmetries in $\varphi$, though in the stationary limit, rigidity theorems suggest that any nonaxisymmetric perturbations are part of the “hair” that a black hole will eventually shed [39, 40]. In regard to assumption (ii), the lack of asymptotic flatness in the Universe has been measured with great precision [41], but any effects from a cosmological horizon on the inner-horizon instability are expected to be negligible, both due to the similarity in the RSET calculations here to those of the Kerr–de Sitter case [17], and because the cosmological constant is currently measured to be $\Lambda\sim 10^{-46}\ M_{\astrosun}^{-2}$ in geometrized units, much too small to have any practical influence. As for assumption (iii), black holes (and therefore near-inner-horizon geometries) form under chaotic, rapidly evolving conditions, and even after settling down, the most quiescent black holes are still nonvacuum and accrete more than enough radiation to trigger the mass inflation instability and destroy any vacuum inner horizon [9]. Additionally, physically relevant fields in the Standard Model are those with spin 1 (electromagnetic), 1/2 (fermionic), or 2 (gravitational), so using a massless scalar field with spin 0 is a simplification seen as a proxy for, e.g., a single photonic degree of freedom. Even regarding assumption (iv), which provides a more general notion than the strict separability that itself is justified by the observation of long-lived accretion disks (lack of separability implies chaotic, destabilizing particle orbits [35]), it will be seen in what follows that conformal separability, while valid in the weak backreaction domain, eventually breaks down once the vierbein of Eq. (16) ceases to be diagonal.
@@
The Einstein tensor corresponding to this spacetime is far too intricate to be presented here in any complete, meaningful way. However, at least in the quasistationary limit (in particular, if the redefined conformal piece $\rho$ is written as $\rho(r,\theta,t)=\text{e}^{\tilde{v}t}\sqrt{r^{2}+a^{2}\cos^{2}\!\theta}$ for some vanishing small accretion parameter $\tilde{v}$), the field equations have already been analyzed in detail in Refs. [5, 6, 7]. The resulting behavior provides classical justification for the decomposition of Eq. (14), since near the inner horizon, the functions $\Delta_{\theta}$, $\Omega_{r}$, and $\Omega_{\theta}$ remain approximately fixed at their Kerr values near the inner horizon, while the horizon function $\Delta_{r}$ and the inflationary function $\xi$ become the dominant contributors to the evolution of the Einstein tensor. The resulting tensor behaves at least initially like a perfect fluid with equal ingoing and outgoing streams, with dominant inflating tetrad-frame components $G_{00}\sim G_{11}\sim G_{01}\sim\Delta_{r}^{-1}$, where the off-diagonal component $G_{01}$ is suppressed by a factor of $\tilde{v}$. Classically, the evolution can then be continued to show that $\Delta_{r}$ eventually stalls out at some tiny value [6], after which $\xi$ grows large and the spacetime undergoes a series of BKL-like bounces toward a strong, spacelike singularity [8].
How might semiclassical effects modify this picture? The tetrad-frame components of the RSET can be reexpressed in terms of the double-null Eddington-Finkelstein coordinates used throughout this paper:
@@
Far enough away from the inner horizon, all components of the RSET should be completely negligible, as they are suppressed by the Planck scale. Thus, everywhere outside the black hole as well as inside when $v$ is not large, a vacuum source should recover the standard Kerr solution with $\xi=0$ and $\Delta_{r}=(r^{2}+a^{2}-2Mr)/(r^{2}+a^{2})^{2}$. However, in the weak backreaction domain, once the null components of the RSET cease to be vanishingly small, Equation (17) indicates that the radial and time components of the tetrad-frame Einstein tensor will begin to diverge as $\Delta_{r}^{-1}$ as one takes $\Delta_{r}\to 0$.
Though the RSET components ${\langle T^{-}_{uv}\rangle}$, ${\langle T^{-}_{u\varphi}\rangle}$, and ${\langle T^{-}_{\varphi\varphi}\rangle}$ have not been computed explicitly, one can make the assumption that their contributions to Eq. (17) will be subdominant compared to that of the double-null components, and more importantly, that they will not contrive to cause any of the specific combinations in the equations above to cancel exactly. Even if they do, then the classical mass inflation phenomenon described in Refs. [5, 6, 7] will take over the evolution, and a spacelike singularity will form. But based on the numerical results in Sec. III, it is apparent that ${\langle T^{-}_{00}\rangle}$ and ${\langle T^{-}_{11}\rangle}$ are the dominant contributors to the evolution in the weak backreaction domain, while the off-diagonal components ${\langle T^{-}_{01}\rangle}$ are also important but initially much smaller than their diagonal counterparts (recall from Fig. 3 that the difference ${\langle T^{-}_{uu}\rangle}-{\langle T^{-}_{vv}\rangle}$ is almost always several orders of magnitude smaller than either individual component).
@@
$-2\partial_{r_{*}}\ln\left(\frac{\rho}{1-\Omega_{r}\Omega_{\theta}}\right)\partial_{r_{*}}\xi\bigg{)}+F_{rt},$ (18)
where the tortoise coordinate $r_{*}$ is defined by $dr/dr_{*}=\text{e}^{3\xi}R^{2}\Delta_{r}$, and the function $F_{rt}$ encompasses subdominant terms related to the precise nature of any classical accretion contributing to nonstationary conformal dynamics.
The dominant term in Eq. (IV.3) is the one involving the second derivative of the inflationary exponent $\xi$ [6]. Equating this term to the semiclassical source from Eq. (17), which behaves as $\Delta_{r}^{-1}$ near the inner horizon, yields after integration the approximate solution
@@
$-2\Omega_{r}(\langle T^{-}_{u\varphi}\rangle-\langle T^{-}_{v\varphi}\rangle)+\Omega_{r}^{2}\langle T^{-}_{\varphi\varphi}\rangle\Big{)},$ (19)
as long as $\xi$ remains smaller than its derivatives. In the regime where the double-null components of the RSET dominate over the shearing components, Eq. (IV.3) thus predicts that when the sum ${\langle T^{-}_{uu}\rangle+\langle T^{-}_{vv}\rangle}$ is positive (negative), the spacetime’s conformal factor $\text{e}^{-\xi}$ will abruptly contract (expand) as the inner horizon is approached at $\Delta_{r}\to 0_{-}$. These initial tendencies align precisely with that of the spherically symmetric case, since ${\langle T^{-}_{uu}\rangle+\langle T^{-}_{vv}\rangle}$ is usually the same sign as ${\langle T^{-}_{vv}\rangle}$.
One can continue the evolution of the spacetime by examining the behavior of the dynamical vierbein of Eqs. (16a)–(16d). Under the assumption that the fixed vierbein of Eqs. (15a)–(15d) remains stable as the initial backreaction ignites the inflationary tendencies described above, the geometry will be governed solely by a diagonal, dynamical, homogeneous line element sourced by inflating streams of semiclassical matter. The classical counterpart of this analysis has already been carried out in detail in Ref. [9], yielding the so-called inflationary Kasner metric: as one might anticipate, the streams of matter are amplified in energy density as $\Delta_{r}$ plunges to zero and $\xi$ increases rapidly. Then, the spacetime undergoes a bounce as $\Delta_{r}$ freezes out at a small value and the $\text{e}^{\xi/2}$ term in Eq. (16b) takes over and causes a radial expansion. The solution works for both positive and negative sources of stress energy, generically producing a series of inflating Kasner-like bounces toward a spacelike singularity.